∮[(z - 1)/(z(z - t)(z - 3i)) dz] ; |z - i| = 1/2


1
Expert's answer
2021-05-12T01:52:12-0400

"\oint f(z)/(z-a) dz =2\pi i f(a)"

"\oint [(z-1)/z(z-t) (z-3i)] dz"

"\u222e[(z\u22121)\/(z^2 \n \u2212tz)(z\u22123i)]dz"

"=\oint [( (z-1)/(z^2-tz))/z-3i] dz = 2\pi f(a)""=2\pi i[ ((t-1)ln(|z-t|)) /(t^2-3it)+((1-3i)ln(|z-3i |)) /(3it+9) +((i) ln(|z|))/3t]"